OpenStudy (anonymous):

Which expressions show how 6 • 45 can be rewritten using the distributive property? Choose all answers that are correct. A. 6 • 40 + 6 • 5 B. 6 • 40 − 6 • 5 C. 6 • 50 – 6 • 5 D. 6 • 50 + 5

OpenStudy (anonymous):

Ok first thing is to distribute

OpenStudy (anonymous):

5(9 - 3) becomes what?

OpenStudy (anonymous):

9-3 =6

OpenStudy (anonymous):

No not quite. it becomes 45 - 15

OpenStudy (anonymous):

Ok so now we have 45 - 15 = w • 9 – 5 • 3 no we simplify the right side of the equal sign. Whats 9-5 x 3?

OpenStudy (anonymous):

why would it be 45- 15?

OpenStudy (anonymous):

9-5 x 3 = 12

OpenStudy (anonymous):

because you are distributing the 5 in the set of numbers. 5(9 - 3) is the same as 5 x 9 - 5 x 3

OpenStudy (anonymous):

Good its 12. So now we have 45 - 15 = w x 12. So lets work from left to right. What is 45 -15?

OpenStudy (anonymous):

45-15 = 30

OpenStudy (anonymous):

Awesome! So now we have 30 = w x 12. So laslty we are going to do the opposite of one side to the other. Instead of multiplying by 12 on the right, lets divide by 12 on the left. SO- 30/12 = w What is 30/12?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

Awesome job! So now your final answer is~ 2.5 = w Did you understand the process?

OpenStudy (anonymous):

hopefully I didnt confuse you!

OpenStudy (anonymous):

yea its a little hard tho BUT i will get better

OpenStudy (anonymous):

Awesome. I am GREAT with these kind of problems. So add me and tag me into any of these anytime if you need help! ok? Good luck

OpenStudy (anonymous):

ok and thx

OpenStudy (anonymous):

dont go yet

OpenStudy (anonymous):

Which expressions show how 6 • 45 can be rewritten using the distributive property? Choose all answers that are correct. A. 6 • 40 + 6 • 5 B. 6 • 40 − 6 • 5 C. 6 • 50 – 6 • 5 D. 6 • 50 + 5